Calculate the solubility of nickel carbonate, NiCO3 in units of grams per liter. Ksp(NiCO3) = 6.6E-9. solubility = g/L
Answer #1

Let molax solubility is S. Ni coz - Ni+2+ Co S s Ksp = [Ni +2] [C032] 6-6x109 – (S) (S) s² = 6-6x1019 S = V6-6x169 molel! S =